OpenStudy (anonymous):

i need help with exponents and fractions the problem is ( 1 over -3)-3. can someone help

OpenStudy (saifoo.khan):

\[(\frac13)^{-3} = 3^3 = 27\]

OpenStudy (anonymous):

im sorry its (-1 over positive 3)-3 but thank you so much that was really fast

OpenStudy (anonymous):

but that was a great answer for that one

OpenStudy (saifoo.khan):

You welcome! =]

OpenStudy (lgbasallote):

-10/3

OpenStudy (anonymous):

ty

OpenStudy (lgbasallote):

that's -1/3 - 3 which is -1/3 - 9/3 = -10/3

OpenStudy (saifoo.khan):